# Generating, Reconstructing, and Representing Discrete and Continuous Data: Generalized Diffusion with Learnable Encoding-Decoding

### v1

@@ Abstract
The vast applications of deep generative models are anchored in three core capabilities—generating new instances, reconstructing inputs, and learning compact representations—across various data types, such as discrete text/protein sequences and continuous images. Existing model families, like Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), autoregressive models, and diffusion models, generally excel in specific capabilities and data types but fall short in others. We introduce generalized diffusion with learnable encoder-decoder (DiLED), that seamlessly integrates the core capabilities for broad applicability and enhanced performance. DiLED generalizes the Gaussian noising-denoising in standard diffusion by introducing parameterized encoding-decoding. Crucially, DiLED is compatible with the well-established diffusion model objective and training recipes, allowing effective learning of the encoder-decoder parameters jointly with diffusion. By choosing appropriate encoder/decoder (e.g., large language models), DiLED naturally applies to different data types. Extensive experiments on text, proteins, and images demonstrate DiLED’s flexibility to handle diverse data and tasks and its strong improvement over various existing models.
@@ 1 Introduction
Figure 1: (a) Generation, reconstruction, and representation are the core capabilities for diverse applications. (b) DiLED shows comprehensive abilities on different text tasks in the customer-review domain.
@@
Existing deep generative models typically show strengths in some, but not all, of the three capabilities, resulting in limited applicability or suboptimal performance (Figure 2). For example, Variational Autoencoders (VAEs, Kingma & Welling, 2014) are known for their tradeoff between realistic generation and faithful reconstruction (Chen et al., 2017; Higgins et al., 2017), especially on text sequences (Bowman et al., 2016; Yang et al., 2017; Li et al., 2020; Liu et al., 2023a); Generative Adversarial Networks (GANs, Goodfellow et al., 2014) inherently lack inference of latent representations. Despite the rich subsequent research of incorporating latent inference (Zhou et al., 2019; Xia et al., 2023a; Donahue et al., 2016; Dumoulin et al., 2017; Li et al., 2017) such as GAN inversion (Xia et al., 2023b; Karras et al., 2020b), they fall short of faithfully reconstructing the inputs (Preechakul et al., 2022); Autoregressive models (e.g., modern large language models) excel in generating text, but often with limited diversity and lacking compact semantic representations(Brown et al., 2020). Similarly, recent diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) deliver new state-of-the-arts on synthesizing photo-realistic images, yet lack compact data representation and often rely on separately trained components for remediation (Rombach et al., 2021; Preechakul et al., 2022).
@@
In this paper, we investigate a new deep generative approach based on generalized diffusion, that inherently integrates the three core capabilities and offers added flexibility to model both discrete and continuous data. Specifically, starting with the popular variational inference perspective of diffusion models (e.g., DDPM, Ho et al., 2020), we show that the standard formulation can be generalized by plugging in arbitrary “noising” operation at early diffusion steps, as long as the inverse “denoising” operation is known. We can thus go beyond the common Gaussian noise (or other predefined image degradation, Bansal et al., 2022) and introduce parameterized encoder-decoder in place of noising-denoising. Crucially, the encoder-decoder parameters can naturally be learned together with other diffusion parameters using the original DDPM loss which is well-established for stable and scalable training (Rombach et al., 2021).
@@
More specifically, we introduce encoder-decoder at the first diffusion step. That is, the diffusion process first encodes the input data into a low-dimensional latent vector, followed by common Gaussian noising steps. The resulting approach, DiLED (Diffusion with Learnable Encoding-Decoding), combines a number of desirable properties and overcomes difficulties in aforementioned deep generative models:
@@
(1) Like VAEs and latent diffusion models (which are based on VAEs), DiLED offers compact semantic representation of data. Moreover, thanks to the unified learning of the encoder-decoder within the standard diffusion framework, DiLED obtains a much improved representation space than VAEs and latent diffusion, as shown in §4.
@@
(2) The improved representation space also allows DiLED to avoid the trade-off between generation and reconstruction capabilities observed in VAEs and augmented GANs (Creswell & Bharath, 2019). Therefore, DiLED seamlessly integrates the three core capabilities, leading to more diverse applications and enhanced performance.
@@
(3) The flexibility of the generalized diffusion formulation allows us to specify any desired encoder-decoder for modeling both discrete and continuous data. DiLED thus overcomes the difficulty of standard diffusion and GANs on text and other discrete modalities.
@@
(4) Moreover, as described in §4, we could further plug in large pretrained (autoregressive) language models (LMs) for initializing the encoder-decoder. This leads to greatly improved performance than previous text diffusion models (Li et al., 2022; Lin et al., 2023; Wu et al., 2023) that are inherently incompatible with off-the-shelf pretrained LMs.
@@
We conduct extensive experiments on text, images, and protein sequences. DiLED demonstrates comprehensive capabilities across a wide range of tasks, such as data synthesis, reconstruction, interpolation, editing, and optimization (e.g., Figure 1b) on the different data modalities. The enhanced flexibility and performance of DiLED over the diverse existing generative models shows the strong potential of the new approach for broad applications.
@@ 2 Background
We begin by providing a brief background of the popular diffusion model formulation from the variational inference perspective (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021a). The training objective and model configurations have been well-established and are effective for stable and scalable training in practice (Rombach et al., 2021), showing advantages to alternative diffusion formalisms and other generative models. Our approach (§3) fits seamlessly in the (generalized) formulation and thus inherits these advantages.
@@
Diffusion formulations (Sohl-Dickstein et al., 2015; Ho et al., 2020) introduce a series of derivations to simplify the above objective, by splitting the sum and transforming and cancelling out relevant terms. The lower bound is thus rewritten as:
@@
Despite their strengths, diffusion models have several limitations. The noisy nature of the latent variables ${\mathbf{x}}_{t}$, which represent corrupted versions of the data (${\mathbf{x}}_{0}$), hinders their ability to capture abstract semantic meaning, making them less suitable for applications requiring such representations (Preechakul et al., 2022). Additionally, the reliance on predefined additive Gaussian noise restricts them to continuous, fixed-length data, posing challenges for handling discrete sequences of varying lengths. Recent efforts (Li et al., 2022; Lin et al., 2023; Wu et al., 2023) adapt the approach to text modeling, but their performance still lags behind pre-trained autoregressive models.
3 Generalized Diffusion with Learnable Encoding-Decoding (DiLED)
We describe DiLED, which integrates the core capabilities of generation, reconstruction, and compact representation within one framework. DiLED thus combines the various advantages and applicabilities of different existing deep generative models while overcoming their limitations.
DiLED generalizes the standard diffusion formulation in §2. The intuition is that the leading steps ($t=0,1,\dots$) in the diffusion process can go beyond using only the common Gaussian noising-denoising. Instead, we can plug in more sophisticated “noising” and “denoising” operations. In particular, we use neural encoder and decoder with learnable parameters in place of “noising” and “denoising”, respectively. The encoder allows the model to map its inputs into a low-dimensional space for a compact representation, and the decoder is to invert the encoding step and recover the inputs. Crucially, this change is fully compatible with the diffusion objective in Eqs. (3) and (4), allowing us to retain most of the critical derivations in the well-established formulation (Sohl-Dickstein et al., 2015; Ho et al., 2020) and thus inherit the stable effective training of both the original denoising parameters and the new encoder-decoder parameters, this is unlike VAEs where the trade-off between KL-divergence and other terms is hard to handle, or LSGM where they make approximations in each component based on certain assumptions. These may all lead to instability during training.
@@ 3.1 The Generalized Formulation
We introduce learnable encoder-decoder to the leading $n$ steps in the diffusion process (i.e., for all $1\leq t\leq n$). For simplicity and clarity, this work considers only the first step (i.e., $t=n=1$) in the diffusion process, though all the derivation can be applied to larger $n$ for $n$-layer hierarchical representations.
@@
As mentioned earlier, we can generalize the standard diffusion objective (Eqs. 3 and 4) to plug in the encoder and decoder seamlessly. The full derivations are provided in the appendix (§A). Briefly, we can adapt the first equation of Eq. (4) as:
@@
where $\rho:=\frac{\alpha_{1}(1-\bar{\alpha}_{1})}{\beta_{1}^{2}}$ and, following the well-validated approximation in Ho et al. (2020), $\gamma_{1}$ is set to $1$ in practice.
@@
This terms is the data reconstruction loss that corresponds to $\mathcal{L}_{0}$ in standard diffusion objective Eq. (4). However, thanks to the incorporation of the decoder $p_{\phi}({\mathbf{x}}_{0}|{\mathbf{x}}_{1})$, we can generalize the original $\mathcal{L}_{0}$ (which is only for continuous data like image) to model various data modalities, such as text and protein sequences of varying lengths. For instance, for text data, the reconstruction loss becomes a standard sequential cross-entropy loss.
We provide more detailed derivations of the full final objective in §A.4 and the complete training process in §B. Thanks to the seamless alignment with the standard diffusion formulation (Ho et al., 2020)), we are allowed to largely follow their training recipes for effective training. In practice, we introduce a hyperparameter weight $w$ to balance the terms $\mathcal{L}_{t-1}$ / $\mathcal{L}_{\text{align}}$ and $\mathcal{L}_{\text{rec}}$.4). In our experiments, using validation data, we set $w=8$ for text and $w=1$ for both protein and images.
@@ 3.2 Generation, Reconstruction, and Representation
A trained DiLED model can naturally support the three core functionalities including generation, reconstruction, and representation, as well as the diverse applications built on top of the functionalities. To generate new samples, similar to standard diffusion, DiLED simulates a random noise ${\mathbf{x}}_{T}$ from prior $p({\mathbf{x}}_{T})$, and goes through the denoising process followed by decoding into a sample $\mathbf{\hat{x}}_{0}$ with the learned decoder. Due to the stochastic nature of the process, such as drawing random noises, DiLED can generate highly diverse samples (§4.1). Representing an input with a compact vector is done straightforwardly by applying the learned encoder. To reconstruct the input, we further apply the decoder on the latent vector to obtain the reconstructed sample. The compact representation space also facilitates other tasks (Figure 1), such as interpolation which is done by drawing an intermediate point between the representations of two given samples and decoding it into the data space, and editing which modifies the sample representation (e.g., with a latent classifier or latent vector arithmetic, as in §4) followed by decoding. We disucss more details and analysis in our experiments (§4).
@@ 3.3 Connections with Other Generative Models
We discuss the rich connections between DiLED to other diverse deep generative models (Goodfellow et al., 2016; Hu & Xing, 2022), providing insights into the integrated advantages of the new approach. DiLED can be viewed as a VAE (Kingma & Welling, 2014) with a learned diffusion model prior. This can be seen through Eq. (7) where $\mathcal{L}_{\text{rec}}$ corresponds to the reconstruction loss in VAE and $\mathcal{L}_{\text{align}}$ corresponds to the KL regularization with prior. In this perspective, the vanilla VAE uses a standard Gaussian prior $p({\mathbf{x}}_{1})$ while DiLED learns the “prior” $p_{\theta}({\mathbf{x}}_{1}|{\mathbf{x}}_{2})$ through diffusion modeling. Note that previous studies have also explored learnable priors in VAE to overcome the difficulty of vanilla VAE training and to improve expressiveness (Dilokthanakul et al., 2016; Chen et al., 2017; Tomczak & Welling, 2018; Razavi et al., 2019; Wehenkel & Louppe, 2021; Vahdat et al., 2021; Goyal et al., 2017). Wehenkel & Louppe (2021) presented a preliminary study of diffusion priors in VAEs but without full derivations and experiments. If we replace all the noising-denoising steps in diffusion with parameterized encoders-decoders, we arrive at a hierarchical VAE model (e.g., NVAE, HiVAE, HVAE Vahdat & Kautz, 2020b; Liu et al., 2023b; Bai et al., 2023). Compared to hierarchical VAEs, DiLED and diffusion allow for more effective training for generation thanks to the special parameterizations and training recipes (Ho et al., 2020). Latent diffusion (Esser et al., 2021) combines VAE with diffusion but trains them separately. DiLED, inherently with unified training, offers a better semantic representation space than latent diffusion whose representation space is that of VAE. Diffusion Autoencoders (DiffAE, Preechakul et al., 2022) and LSGM (Vahdat et al., 2021) explore joining training strategies of diffusion and autoencoder. DiLED, benefiting from the well-established training objective and recipe from variational diffusion (Ho et al., 2020), achieves more effective training and better performance (§4.2). Also, the applications of DiffAE and LSGM on text/protein data have not been explored. Compared to recent text diffusion models (Yuan et al., 2023; He et al., 2022; Lin et al., 2023; Ye et al., 2023) that generate text non-autoregressively, DiLED can flexibly accommodate pretrained autogressive language models as the decoder for much enhanced performance.

### v2

# Unified Generation, Reconstruction, and Representation: Generalized Diffusion with Adaptive Latent Encoding-Decoding

@@ Abstract
The vast applications of deep generative models are anchored in three core capabilities—generating new instances, reconstructing inputs, and learning compact representations—across various data types, such as discrete text/protein sequences and continuous images. Existing model families, like variational autoencoders (VAEs), generative adversarial networks (GANs), autoregressive models, and (latent) diffusion models, generally excel in specific capabilities and data types but fall short in others. We introduce Generalized Encoding-Decoding Diffusion Probabilistic Models (EDDPMs) which integrate the core capabilities for broad applicability and enhanced performance. EDDPMs generalize the Gaussian noising-denoising in standard diffusion by introducing parameterized encoding-decoding. Crucially, EDDPMs are compatible with the well-established diffusion model objective and training recipes, allowing effective learning of the encoder-decoder parameters jointly with diffusion. By choosing appropriate encoder/decoder (e.g., large language models), EDDPMs naturally apply to different data types. Extensive experiments on text, proteins, and images demonstrate the flexibility to handle diverse data and tasks and the strong improvement over various existing models. Code is available at https://github.com/guangyliu/EDDPM
@@
Figure 1: (a) Generation, reconstruction, and representation are the core capabilities for diverse applications. (b) EDDPM shows comprehensive abilities on different text tasks in the customer-review domain (§4.1).
@@ 1 Introduction
Existing deep generative models typically show strengths in some, but not all, of the three capabilities, resulting in limited applicability or suboptimal performance (Figure 2). For example, variational autoencoders (VAEs, Kingma & Welling, 2014) are known for their tradeoff between realistic generation and faithful reconstruction (Chen et al., 2017; Higgins et al., 2017a), especially on text sequences (Bowman et al., 2016; Yang et al., 2017; Li et al., 2020; Liu et al., 2022, 2023a); Generative adversarial networks (GANs, Goodfellow et al., 2014) inherently lack inference of latent representations. Despite the rich subsequent research of incorporating latent inference (Zhou et al., 2019; Xia et al., 2023a; Donahue et al., 2016; Dumoulin et al., 2017; Li et al., 2017) such as GAN inversion (Xia et al., 2023b; Karras et al., 2020b), they fall short of faithfully reconstructing the inputs (Preechakul et al., 2022); Autoregressive models (e.g., modern large language models) excel in generating text, but often with limited diversity and lacking compact semantic representations (Brown et al., 2020). Similarly, recent diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) deliver new state-of-the-arts on synthesizing photo-realistic images, yet lack compact data representation and often rely on separately trained components for remediation (Rombach et al., 2021; Preechakul et al., 2022).
@@
In this paper, we investigate a new deep generative approach based on generalized diffusion, that inherently integrates the three core capabilities and offers added flexibility to model both discrete and continuous data. Specifically, starting with the popular variational inference perspective of diffusion models (as in DDPM, Ho et al., 2020), we show that the standard formulation can be generalized by plugging in arbitrary “noising” operation at early diffusion steps, as long as the inverse “denoising” operation can be modeled. We can thus go beyond the common Gaussian noise (or other predefined image degradation, Bansal et al., 2022) and introduce parameterized encoder-decoder in place of noising-denoising. Crucially, the encoder-decoder parameters can naturally be learned together with other diffusion parameters using the original DDPM framework which is well-established for stable and scalable training (Rombach et al., 2021).
@@
More specifically, we introduce encoder-decoder at the first diffusion step. That is, the diffusion process first encodes the input data into a low-dimensional latent vector, followed by common Gaussian noising steps. The resulting approach, EDDPMs (Generalized Encoding-Decoding Dffusion Probabilistic Models), combine a number of desirable properties and overcome difficulties in aforementioned deep generative models:
  1. (1)
@@
    Like VAEs and latent diffusion models (which are based on VAEs), EDDPMs offer compact semantic representation of data. Moreover, thanks to the unified learning of the encoder-decoder within the standard diffusion framework, EDDPMs obtain a much improved representation space than VAEs and latent diffusion, as shown in §4.
  2. (2)
@@
    The improved representation space also allows EDDPMs to avoid the trade-off between generation and reconstruction capabilities observed in VAEs and augmented GANs (Creswell & Bharath, 2019). Therefore, EDDPMs seamlessly integrate the three core capabilities, leading to more diverse applications and enhanced performance.
  3. (3)
@@
    The flexibility of the generalized diffusion formulation allows us to specify any desired encoder-decoder for modeling both discrete and continuous data. EDDPMs thus overcome the difficulty of standard diffusion and GANs on text and other discrete modalities.
  4. (4)
@@
    Moreover, as described in §4, we could further plug in large pretrained (autoregressive) language models (LMs) for initializing the encoder-decoder. This leads to greatly improved performance than previous text diffusion models (Li et al., 2022; Lin et al., 2023; Wu et al., 2023) that are inherently incompatible with off-the-shelf pretrained LMs.
@@
We conduct extensive experiments on text, images, and protein sequences. EDDPMs demonstrate comprehensive capabilities across a wide range of tasks, such as data synthesis, reconstruction, interpolation, editing, and optimization (e.g., Figure 1b) on the different data modalities.
The unified capabilities and enhanced performance of EDDPMs compared to existing generative models demonstrate its significant potential as a foundational technique for developing new, broadly applicable foundation models.
@@ 2 Background
We first give brief background of the popular diffusion model formulation from the variational inference perspective (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021a). The training objective and model configurations have been well-established and are effective for stable and scalable training in practice (Rombach et al., 2021), showing advantages to alternative diffusion formalisms and other generative models. Our approach (§3) fits seamlessly in the (generalized) formulation and inherits these advantages.
@@
Recent diffusion formulation, in particular the denoising diffusion probabilistic models (DDPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020), introduces a series of derivations to simplify the above objective, by splitting the sum and transforming and cancelling out relevant terms. The lower bound is thus rewritten as:
@@
Despite their strengths, diffusion models have several limitations. The noisy nature of the latent variables ${\mathbf{x}}_{t}$, which represent corrupted versions of the data (${\mathbf{x}}_{0}$), hinders their ability to capture abstract semantic meaning, making them less suitable for applications requiring such representations (Preechakul et al., 2022). Additionally, the reliance on predefined additive Gaussian noise restricts them to continuous, fixed-length data, posing challenges for handling discrete sequences of varying lengths. Recent efforts (Li et al., 2022; Lin et al., 2023; Wu et al., 2023) adapt the approach to text modeling, but their performance still greatly lags behind pretrained autoregressive models.
3 Generalized Encoding-Decoding Diffusion Probabilistic Models (EDDPMs)
We describe EDDPMs, which integrate the core capabilities of generation, reconstruction, and compact representation within one framework. EDDPMs thus combine the various advantages and applicabilities of different existing deep generative models while overcoming their limitations.
EDDPMs generalize the standard diffusion formulation described in §2. The intuition is that the leading steps ($t=0,1,\dots$) in the diffusion process can go beyond using only the common Gaussian noising-denoising. Instead, we can plug in more sophisticated “noising” and “denoising” operations. In particular, we use neural encoder and decoder with learnable parameters in place of “noising” and “denoising”, respectively. The encoder allows the model to map its inputs into a low-dimensional space for a compact representation, and the decoder is to invert the encoding step and recover the inputs. Crucially, this change is fully compatible with the diffusion objective in Eqs. (3) and (4), allowing us to retain most of the critical derivations in the well-established formulation (Sohl-Dickstein et al., 2015; Ho et al., 2020) and thus inherit the stable effective training of both the original denoising parameters and the new encoder-decoder parameters. As we discussed in §3.3, this differs crucially from latent diffusion models that train a VAE and a diffusion separately and are limited by the capability and quality of the VAE component. Our approach also shows superiority over previous joint training (Preechakul et al., 2022; Vahdat et al., 2021) which requires various approximations and can lead to unstable training and inferior performance.
@@ 3.1 The Generalized Formulation
We introduce learnable encoder-decoder to the leading $n$ steps in the diffusion process (i.e., for all steps $t$, where $1\leq t\leq n$). For simplicity and clarity, this work considers only the first step (i.e., $t=n=1$) in the diffusion process, though all the derivation can be applied to larger $n$ for $n$-layer hierarchical representations.
@@
As mentioned earlier, we can generalize the standard diffusion objective (Eqs. 3 and 4) to plug in the encoder and decoder seamlessly. The full derivations are provided in the appendix (§A). Briefly, we can adapt the left-hand side of Eq. (4) as:
@@
where ${\bm{\mu}}_{\theta}$ is the Gaussian mean in Eq. (2); $\rho:=\frac{\alpha_{1}(1-\bar{\alpha}_{1})}{\beta_{1}^{2}}$ and, following the approximation in Ho et al. (2020), $\gamma_{1}$ is set to $1$ in practice.
@@
This terms is the data reconstruction loss that corresponds to $\mathcal{L}_{0}$ in standard diffusion objective Eq. (4). However, thanks to the incorporation of the decoder $p_{\phi}({\mathbf{x}}_{0}|{\mathbf{x}}_{1})$, we can generalize the original $\mathcal{L}_{0}$ (which is only for continuous data like image) to model various data modalities, such as text and protein sequences of varying lengths. For instance, for text data, the decoder can be a (pretrained) language model and the reconstruction loss becomes a standard sequential cross-entropy loss.
We provide more detailed derivations of the full final objective in §A.4 and the complete training process in §B. Thanks to the compatibility with the standard diffusion formulation (Ho et al., 2020)), we can largely follow their training recipes for effective training. In practice, we introduce a hyperparameter weight $w$ to balance the term $\mathcal{L}_{\text{rec}}$ against other terms like $\mathcal{L}_{t-1}$ and $\mathcal{L}_{\text{align}}$, as detailed in §A.4. In our experiments, using validation data, we set $w=8$ for text and $w=1$ for both protein and images.
@@ 3.2 Generation, Reconstruction, and Representation
The trained EDDPMs can naturally support the three core functionalities including generation, reconstruction, and representation, as well as the diverse applications built on top of the functionalities. To generate new samples, similar to standard diffusion, EDDPMs simulate a random noise ${\mathbf{x}}_{T}$ from prior $p({\mathbf{x}}_{T})$, and go through the denoising process followed by decoding into a sample $\mathbf{\hat{x}}_{0}$ with the learned decoder. Due to the stochastic nature of the process (e.g., drawing random noise ${\mathbf{x}}_{T}$), EDDPMs can generate highly diverse samples (§4.1). Representing an input with a compact vector is done straightforwardly by applying the learned encoder. To reconstruct the input, we further apply the decoder on the latent vector to obtain the reconstructed sample. The compact representation space also facilitates other tasks (Figure 1), such as interpolation which is done by drawing an intermediate point between the representations of two given samples and decoding it into the data space, and editing which modifies the sample representation (e.g., with a latent classifier or latent vector arithmetic) followed by decoding. We disucss more details and analysis in experiments (§4).
@@ 3.3 Connections with Other Generative Models
We discuss the rich connections between EDDPMs to other diverse deep generative models (Goodfellow et al., 2016; Hu et al., 2018; Hu & Xing, 2022), providing insights into the integrated advantages of the new approach. EDDPMs can be viewed as VAEs with a learned diffusion model prior. This can be seen through Eq. (7) where $\mathcal{L}_{\text{rec}}$ corresponds to the reconstruction loss in VAEs and $\mathcal{L}_{\text{align}}$ corresponds to the KL regularization with prior. In this perspective, the vanilla VAEs use a standard Gaussian prior $p({\mathbf{x}}_{1})$ while EDDPMs learn the “prior” $p_{\theta}({\mathbf{x}}_{1}|{\mathbf{x}}_{2})$ through diffusion modeling. Note that previous studies have also explored learnable priors in VAEs to overcome the difficulty of vanilla VAEs training and to improve expressiveness (Dilokthanakul et al., 2016; Chen et al., 2017; Tomczak & Welling, 2018; Razavi et al., 2019; Wehenkel & Louppe, 2021; Vahdat et al., 2021; Goyal et al., 2017). Wehenkel & Louppe (2021) presented a preliminary study of diffusion priors in VAEs but without full derivations and experiments. If we replace all the noising-denoising steps in diffusion with parameterized encoders-decoders, we arrive at a hierarchical VAEs (e.g., NVAE, HiVAE, HVAE, Vahdat & Kautz, 2020b; Liu et al., 2023b; Bai et al., 2023). Compared to hierarchical VAEs, EDDPMs and diffusion allow for more effective training for generation thanks to the special parameterizations and training recipes (Ho et al., 2020). Latent diffusion (Esser et al., 2021) combines VAEs with diffusion but trains them separately. EDDPMs, inherently with unified training, offer a better semantic representation space than latent diffusion whose representation space is that of the vanilla VAEs. Diffusion Autoencoders (DiffAE, Preechakul et al., 2022) and LSGM (Vahdat et al., 2021) explore joining training strategies of diffusion and autoencoder and introduce different approximation for tractability. EDDPMs, benefiting from the well-established training objective and recipe from variational diffusion (Ho et al., 2020), achieve more effective training and better performance (§4.2). Also, the applications of DiffAE and LSGM on text/protein data have not been explored. Compared to recent text diffusion models (Yuan et al., 2023; He et al., 2022; Lin et al., 2023; Ye et al., 2023) that generate text non-autoregressively, EDDPMs can flexibly accommodate pretrained autogressive language models as the decoder for much enhanced performance.